OpenStudy (loser66):

Riemann theorem: Let f(x) =x^2, and let P denote a partition of an interval [a,b] i) Let L(f,P), M(f,P)and R(f,P) denote respectively the left end point, the midpoint and the right endpoint sums of f(x) on P. Show that L+4M+R = 2(b^3-a^3). I got this part ii) Show that for any partition P with \(||P||<\dfrac{\varepsilon}{b^2-a^2}, R-L<\varepsilon\). Please, help

OpenStudy (loser66):

@freckles

OpenStudy (loser66):

My attempt: \(R(f,P, x_i) = \sum_{i=1}^n f(x_i)||P||\) \(L(f,P, x_{i-1})= \sum_{i=1}^n f(x_{i-1})||P||\) -------------------------------------------- \(R-L = \sum_{i=1}^n(f(x_i)-f(x_{i-1})||P||\) Let \(\delta =\dfrac{\varepsilon}{b^2-a^2}\), then \(||P||<\delta\)

OpenStudy (loser66):

and \((R-L) <\delta \sum_{i=1}^n (x_i^2-x_{i-1}^2)=\varepsilon\) on [a,b]

OpenStudy (loser66):

iii) from i) and ii) conclude that for any P with \(||P||< \dfrac{\varepsilon}{b^2-a^2}\) and for any choice of test points \(x_i^*, x_i^#\in [x_{i-1}, x_i] |R(f,P,x_i^*)-R(f,P,x_i^#)|<\varepsilon\)

OpenStudy (loser66):

I doubt myself for this part, since it is trivial. My argument: if we pick any point in \([x_{i-1},x_i]\), we make a "finerment" partition. Hence for the 2 new points , we have 2 news intervals. The ||P|| works well for the new one. What are we supposed to do then?